OpenStudy (anonymous):

Prove each identity? cos (x+pi/4) + cos (x-pi/4) = sqrt2 cosx

OpenStudy (campbell_st):

use the sum and difference for cos cos(A + B) = cos(A)cos(B) - sin(A)sin(B) and cos(A - B) = cos(A)cos(B) + sin(A)sin(B) you will need to use the exact value of pi/4 for sin and cos... then simplify. and you may also need to rationalise a denominator...

OpenStudy (anonymous):

@campbell_st, I did as you said. However, I am a bit confused on how I would need to rationalize a denominator. Since combining like terms, I ended up with cos sqrt2/2 + cos sqrt2/2. That is the part where I am stuck on. I am not sure if I did something wrong.

OpenStudy (campbell_st):

ok so you get \[\cos(x)\cos(\frac{\pi}{4}) - \sin(x)\sin(\frac{\pi}{4}) + \cos(x)\cos(\frac{\pi}{4}) - \sin(x) \sin(\frac{\pi}{4})\] the terms in sin cancel each other... leaving \[2\cos(x)\cos(\frac{\pi}{4})\] find the exact value of cos(pi/4) and substitute

OpenStudy (campbell_st):

oops there should be a + instead of a - before the 2nd sin(x)sin(pi/4)

OpenStudy (anonymous):

@campbell_st, okay, I got it. I should have combined like terms first instead of distributing. Thank you!